OpenStudy (anonymous):

what is the simplified form of y^2-y-12 overY^2+8y+15

OpenStudy (anonymous):

(y-4)(y+3)/(y+5)(y+3) (y-4)/(y+5)

OpenStudy (anonymous):

\[\left[\begin{matrix}y^2-y- & 12 \\ y^2=+8y & +15\end{matrix}\right]\]

OpenStudy (anonymous):

(y-4)(y+3)/(y+3)(y+5) =y-4 over y+5

OpenStudy (anonymous):

factor the top and bottom then cancel the y+3